# Roberto Dale,22footnotetext: Corresponding author.    Jesús M. Gandía    Juan Antonio Morales-Lladosa    and Ramon Lapiedra

### v1

@@ 7 Violation of the CHSH Inequalities
From the calculations of $\left|\left\langle C\right\rangle\right|$ summarised in Tables 2 (WMAP) and 3 (Planck), several noteworthy observations can be made: (i) all values violate Bell’s inequalities; (ii) the most closely matched values correspond to the W band of WMAP and the NILC pipeline of Planck, with the latter exhibiting the highest value; and (iii) the angular ranges where these violations occur are remarkably similar, spanning $\left[{45^{\circ},51^{\circ}}\right]\cup\left[{129^{\circ},141^{\circ}}\right]$ for WMAP and $\left[{39^{\circ},49^{\circ}}\right]\cup\left[{131^{\circ},141^{\circ}}\right]$ for Planck.
@@
Figure 4 provides a summary of the comparison between the expected values $\alpha\right>$ derived from the degraded maps and the original map. This comparison is presented for the case of the WMAP Q band.
Figure 4: This figure summarises the expected values of $\alpha\right>$ from the WMAP Q band maps, displayed across three panels, comparing the downgraded resolution ($r=6$) with the original resolution ($r=9$). The left panel shows four curves: the dash-dotted black line represents the map downgraded to $r=6$ using a neighbour averaging method, the dashed green line represents the map downgraded to $r=6$ using spherical harmonics, and the two red lines correspond to the original WMAP data (at $r=9$ resolution) for the intervals $\left[36^{\circ},54^{\circ}\right]$ and $\left[126^{\circ},146^{\circ}\right]$, respectively. The two right panels illustrate the relative differences between each of the $r=6$ resolution maps and the original $r=9$ map, with the top panel corresponding to the interval $\left[36^{\circ},54^{\circ}\right]$ and the bottom panel corresponding to the interval $\left[126^{\circ},146^{\circ}\right]$.
To assess the differences in the expected valuess $\alpha\right>$, we define the function
$\Delta\left(\alpha\right)\equiv\left|{\frac{{{{\left\langle\alpha\right\rangle}_{r9}}-{{\left\langle\alpha\right\rangle}_{r6}}}}{{{{\left\langle\alpha\right\rangle}_{r9}}}}}\right|,$ (7.1)
which provides an indication of the degree of deviation between the degraded maps and the original ones. The terms ${\left\langle\alpha\right\rangle}_{r9}$ and ${\left\langle\alpha\right\rangle}_{r6}$ denote the expected values $\alpha\right>$ for the original and downgraded maps, respectively.
@@
134.50 46.25 135.50 136.25 2.44335 0.00004
134.50 46.75 135.50 136.75 2.44322 0.00004
134.25 46.25 135.75 136.25 2.44319 0.00004
134.50 47.00 135.50 137.00 2.44312 0.00004
134.00 46.25 136.00 136.25 2.44310 0.00004
134.25 46.75 135.75 136.75 2.44307 0.00004
134.75 47.50 136.25 136.50 2.44305 0.00004
134.25 45.25 134.75 136.25 2.44303 0.00004
134.25 45.50 134.75 136.50 2.44300 0.00004
134.00 46.75 136.00 136.75 2.44298 0.00004
In light of the results obtained from the original maps, we can draw the following conclusions: (i) As observed to varying degrees across all degraded maps analysed, it is possible to identify multiple tetrads that violate the cosmic inequalities (2.1); (ii) The maximum values of $\left|{\left\langle C\right\rangle}\right|$ slightly exceed those calculated for the corresponding degraded maps; and (iii) The angle values of each tetrad exhibit remarkable consistency across all cases.
@@ 8 Discussion on the made assumptions and concluding results
The analysis of data from the WMAP and Planck satellites regarding the CMB, using both degraded maps and samples from the original maps, has produced conclusive findings. Even in the most unfavourable scenarios considered, it is possible to identify multiple sets of four sky directions—corresponding to causally disconnected events at the time of decoupling—that violate the cosmic CHSH inequalities (2.1).
@@
The present study is based on intensity data, $I$ (Stokes), provided by the WMAP and Planck satellites. After processing, these data have been utilised to examine the potential violation of the cosmic CHSH inequalities (2.1). As previously mentioned, $Q$ and $U$ (Stokes) polarisation data are also available. In this context, it is noteworthy that our future research efforts in this area will focus on testing Bell-like inequalities involving these polarisation data. To this end, we are currently engaged in both the processing of these data and the derivation of new inequalities.
@@
In summary, the eventual detection of a violation of the cosmic CHSH inequalities cannot be attributed to any potential information transmission connecting the two measurements of the same pair, nor can it stem from information originating from the pre-inflationary universe. Such a violation can only arise from purely quantum correlations.
@@
Figure 5: This figure comprises three panels, each depicting the normalised probability distribution of all possible values of the function $\left|\langle C\rangle\right|$ for the WMAP satellite across its three frequency bands (Q, V, and W). The blue bars represent the results obtained using maps degraded via the neighbour-pixel averaging mechanism, whereas the dashed red step plot corresponds to results derived from maps degraded using spherical harmonics. The dashed black vertical line separates the sets of four directions that violate the CHSH cosmic inequalities (2.1) (on the right) from those that do not (on the left). The common width used for both bars and steps in the representation is $\Delta\left|\langle C\rangle\right|=0.05$.
In the study of CMB maps, the resolution of these maps is typically characterised by the parameter $r$ or the parameter $N_{\rm side}$. To establish a relationship between this characterisation and other properties relevant to our analysis, we present Table 5. The derivation of the mean pixel spacing plays a crucial role in determining the appropriate beam size for the calculation of the expected values $\alpha\right>$. This calculation has been performed using algorithms implemented in the Healpy library [65].
@@
Figure 6: As in Figure 5, the normalised probability distribution of the function $\left|\langle C\rangle\right|$ is presented, this time for the four Planck pipelines, across four panels. In each panel, the blue bars correspond to the neighbour-pixel averaging degradation method, while the dashed red step plot represents the spherical harmonics approach. The vertical dashed black line delineates the cases that violate the inequalities (2.1) (to the right) from those that do not (to the left). The bin width used for both the blue bars and the red step plot is $\Delta\left|\langle C\rangle\right|=0.05$.
While the primary objective of this study is to determine whether the CHSH inequalities (2.1) are violated, and to this end, we have presented a series of tetrads that demonstrate such violations across all WMAP bands and Planck pipelines, additional statistical information may also be of interest to the reader. Accordingly, in this section, we provide a series of figures and tables offering a more comprehensive perspective on the results obtained for the function $\left|\langle C\rangle\right|$.
As anticipated, the distribution of values for the function $\left|\langle C\rangle\right|$ is highly consistent across the three bands of the WMAP satellite, as illustrated in detail in Figure 5. A similar pattern is observed in the percentages of cases where the inequalities (2.1) are violated, as presented in the third and fourth columns of Table 6, corresponding to the condition $\left|\langle C\rangle\right|\geq 2$. Likewise, as expected, the probability distributions obtained from the two degradation methods—blue bars for the neighbour-pixel averaging method and red dashed steps for the spherical harmonics approach—are nearly identical and indistinguishable to the naked eye.
@@
Acknowledgments
This work was partially funded by the Conselleria d’Educació, Universitats i Ocupació of the Generalitat Valenciana through projects CIAICO/2022/252 and ASFAE/2022/005. We gratefully acknowledge the invaluable technical assistance of our colleague Rafael Garbayo from the Computer Administration of the Statistics Department at Universidad Miguel Hernández.

### v2

@@ 7 Violation of the CHSH Inequalities
From the calculations of $\left|\left\langle C\right\rangle\right|$ summarised in Tables 2 (WMAP) and 3 (Planck), several noteworthy observations can be made: (i) all values violate Bell’s inequalities; (ii) the most closely matched values correspond to the W band of WMAP and the NILC pipeline of Planck, with the latter exhibiting the highest value; and (iii) the angular ranges where these violations occur are remarkably similar, spanning $\left[{45^{\circ},49^{\circ}}\right]\cup\left[{131^{\circ},141^{\circ}}\right]$ for WMAP and $\left[{41^{\circ},49^{\circ}}\right]\cup\left[{133^{\circ},141^{\circ}}\right]$ for Planck.
@@
Figure 4 provides a summary of the comparison between the expected values $\left<\alpha\right$> derived from the degraded maps and the original map. This comparison is presented for the case of the WMAP Q band.
Figure 4: This figure summarises the expected values of $\left<\alpha\right$> from the WMAP Q band maps, displayed across three panels, comparing the downgraded resolution ($r=6$) with the original resolution ($r=9$). The left panel shows four curves: the dash-dotted black line represents the map downgraded to $r=6$ using a neighbour averaging method, the dashed green line represents the map downgraded to $r=6$ using spherical harmonics, and the two red lines correspond to the original WMAP data (at $r=9$ resolution) for the intervals $\left[36^{\circ},54^{\circ}\right]$ and $\left[126^{\circ},146^{\circ}\right]$, respectively. The two right panels illustrate the relative differences between each of the $r=6$ resolution maps and the original $r=9$ map, with the top panel corresponding to the interval $\left[36^{\circ},54^{\circ}\right]$ and the bottom panel corresponding to the interval $\left[126^{\circ},146^{\circ}\right]$.
To assess the differences in the expected valuess $\left<\alpha\right$>, we define the function
$\Delta\left(\alpha\right)\equiv\left|{\frac{{{{\left\langle\alpha\right\rangle}_{r9}}-{{\left\langle\alpha\right\rangle}_{r6}}}}{{{{\left\langle\alpha\right\rangle}_{r9}}}}}\right|,$ (12)
which provides an indication of the degree of deviation between the degraded maps and the original ones. The terms ${\left\langle\alpha\right\rangle}_{r9}$ and ${\left\langle\alpha\right\rangle}_{r6}$ denote the expected values $\left<\alpha\right$> for the original and downgraded maps, respectively.
@@
134.75 135.75 136.25 46.75 2,44349 0.00004
134.75 136.00 136.25 47.00 2,44339 0.00004
134.50 135.75 136.25 46.75 2,44335 0.00004
134.75 135.50 136.25 46.50 2,44334 0.00004
134.75 135.25 136.25 46.25 2,44330 0.00004
134.50 136.00 136.25 47.00 2,44325 0.00004
134.75 135.00 136.25 46.00 2,44324 0.00004
134.25 135.50 136.25 45.50 2,44321 0.00004
134.50 135.50 136.25 46.50 2,44320 0.00004
134.50 135.25 136.25 46.25 2,44316 0.00004
In light of the results obtained from the original maps, we can draw the following conclusions: (i) As observed to varying degrees across all degraded maps analysed, it is possible to identify multiple tetrads that violate the cosmic inequalities (1); (ii) The maximum values of $\left|{\left\langle C\right\rangle}\right|$ slightly exceed those calculated for the corresponding degraded maps; and (iii) The angle values of each tetrad exhibit remarkable consistency across all cases.
@@ 8 Discussion on the made assumptions and concluding results
The analysis of data from the WMAP and Planck satellites regarding the CMB, using both degraded maps and samples from the original maps, has produced conclusive findings. Even in the most unfavourable scenarios considered, it is possible to identify multiple sets of four sky directions—corresponding to four pairs of causally disconnected events at the time of decoupling—that violate the cosmic CHSH inequalities (1).
@@
The present study is based on intensity data, $I$ (Stokes), provided by the WMAP and Planck satellites. After processing, these data have been utilised to examine the potential violation of the cosmic CHSH inequalities (1). As previously mentioned, $Q$ and $U$ (Stokes) polarisation data are also available. In this context, it is noteworthy that our future research efforts in this area will focus on testing Bell-like inequalities involving these polarisation data. To this end, we are currently engaged in both the processing of these data and the derivation of new inequalities.
@@
In summary, the eventual detection of a violation of the cosmic CHSH inequalities cannot be attributed to any potential information transmission between the measurements of a given pair, nor to signals originating from the pre-inflationary universe. Such a violation can only arise from purely quantum correlations. In our analysis, these correlations involve temperature fluctuations measured along directions on the sky corresponding to space-like separated regions at the time of last scattering. Assuming a quantum statistical description of the primordial fluctuations—consistent with the standard inflationary framework—such violations provide indirect evidence of entanglement imprinted in the early universe.
@@
In the study of CMB maps, the resolution of these maps is typically characterised by the parameter $r$ or the parameter $N_{\rm side}$. To establish a relationship between this characterisation and other properties relevant to our analysis, we present Table 5. The derivation of the mean pixel spacing plays a crucial role in determining the appropriate beam size for the calculation of the expected values $\left<\alpha\right$>. This calculation has been performed using algorithms implemented in the Healpy library [65].
Figure 5: This figure comprises three panels, each depicting the normalised probability distribution of all possible values of the function $\left|\langle C\rangle\right|$ for the WMAP satellite across its three frequency bands (Q, V, and W). The blue bars represent the results obtained using maps degraded via the neighbour-pixel averaging mechanism, whereas the dashed red step plot corresponds to results derived from maps degraded using spherical harmonics. The dashed black vertical line separates the sets of four directions that violate the CHSH cosmic inequalities (1) (on the right) from those that do not (on the left). The common width used for both bars and steps in the representation is $\Delta\left|\langle C\rangle\right|=0.05$.
@@
Figure 6: As in Figure 5, the normalised probability distribution of the function $\left|\langle C\rangle\right|$ is presented, this time for the four Planck pipelines, across four panels. In each panel, the blue bars correspond to the neighbour-pixel averaging degradation method, while the dashed red step plot represents the spherical harmonics approach. The vertical dashed black line delineates the cases that violate the inequalities (1) (to the right) from those that do not (to the left). The bin width used for both the blue bars and the red step plot is $\Delta\left|\langle C\rangle\right|=0.05$.
While the primary objective of this study is to determine whether the CHSH inequalities (1) are violated, and to this end, we have presented a series of tetrads that demonstrate such violations across all WMAP bands and Planck pipelines, additional statistical information may also be of interest to the reader. Accordingly, in this section, we provide a series of figures and tables offering a more comprehensive perspective on the results obtained for the function $\left|\langle C\rangle\right|$.
As anticipated, the distribution of values for the function $\left|\langle C\rangle\right|$ is highly consistent across the three bands of the WMAP satellite, as illustrated in detail in Figure 5. A similar pattern is observed in the percentages of cases where the inequalities (1) are violated, as presented in the third and fourth columns of Table 6, corresponding to the condition $\left|\langle C\rangle\right|\geq 2$. Likewise, as expected, the probability distributions obtained from the two degradation methods—blue bars for the neighbour-pixel averaging method and red dashed steps for the spherical harmonics approach—are nearly identical and indistinguishable to the naked eye.
@@
Appendix B Unmixed Probabilities: Method and Results
[…]
In this section, however, we summarise the results obtained using the so-called ‘Unmixed Probabilities’ approach. This method is considerably simpler: for each frequency, the sign of $\delta T$ at a given direction determines a deterministic outcome of either +1+1 or 1-1, without accounting for measurement uncertainty. As a consequence, the probabilities entering expression 3 are straightforwardly computed as:
$P(X,Y,\alpha)=\frac{\nu(X,Y,\alpha)}{N_{\alpha}},$ (13)
where $\nu(X,Y,\alpha)$ is the raw count of occurrences of the configuration $(X,Y,\alpha)$, and $N_{\alpha}$ is the total number of such pairings at angular separation $\alpha$.
Figure 7: Comparison between the expected values $\left\langle F(\boldsymbol{\hat{x}})F(\boldsymbol{\hat{x}}),\alpha\right\rangle$ calculated via the Mixed and Unmixed Probabilities methods for representative cases from each satellite mission. The solid black line represents the previously analysed case using the ‘Mixed Probabilities’ method, while the dashed red line corresponds to the new results obtained with the ‘Unmixed Probabilities’ approach. The left panel refers to the WMAP Q-band, while the right one refers to the Planck mission analysed using the COMMANDER pipeline.
In Figure 7, we show two representative cases—one from each mission—to illustrate the differences between the expected values obtained using the ‘Mixed’ and ‘Unmixed Probabilities’ methods. The observed behaviour is consistent across all analysed cases and may be summarised as follows:
  1. (a)
    In the intermediate angular range $\alpha\in[30^{\circ},150^{\circ}]$, there are no appreciable differences between the two methods.
  2. (b)
    However, as $\alpha\to 0^{\circ}$ (or $180^{\circ}$), the results begin to diverge: while the ‘Mixed Probabilities’ curve tends towards +0.90+0.90 (or 0.90-0.90), the Unmixed Probabilities curve asymptotically approaches the ideal values of +1+1 (or 1-1).
This latter result for the ‘Unmixed Probabilities’ case is consistent with expectations. Specifically, as $\alpha\to 0^{\circ}$, the unit vectors $\boldsymbol{\hat{x}}$ and $\boldsymbol{\hat{y}}$ become indistinguishable, and the limiting behaviour $F(\boldsymbol{\hat{x}})F(\boldsymbol{\hat{x}})\to F^{2}(\boldsymbol{\hat{x}})=1$ is indeed fulfilled.
From the results obtained in the calculation of the expected values discussed above, we employed the same code to scan and identify the tetrads of angles $\left\{\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}\right\}$ that maximise the value of the function $\left|\left\langle C\right\rangle\right|$. In Table 7, we present the three highest values obtained using the Q band (WMAP mission) and the COMMANDER pipeline (Planck mission), respectively.
In light of the results presented in Table 7, the following remarks are in order:
  1. (a)
    It can be affirmed that, for both the WMAP Q-band and the Planck COMMANDER pipeline, the maximum values of the function $\left|\left\langle C\right\rangle\right|$ increase by approximately $1.5\%$ compared to those previously obtained using the ‘Mixed Probabilities’ method, as presented in Tables 2 and 3. This leads us to conclude that the primary scenario based on the ‘Mixed Probabilities’ method is indeed the most conservative with respect to our claims regarding the potential violation of cosmic Bell inequalities.
  2. (b)
    The sets of angle tetrads shown in Table 7 are identical to those previously obtained using the ‘Mixed Probabilities’ method and reported in Tables 2 and 3, for both the WMAP Q-band and the Planck COMMANDER data analyses.
  3. (c)
    Lastly, although the ordering of these tetrads remains the same in the WMAP case, the second and third highest values are interchanged in the Planck results. However, it is important to note that the corresponding $\left|\left\langle C\right\rangle\right|$ values in these two cases are extremely close, practically indistinguishable between both probability calculation methodologies.
Table 7: The three highest values of the function $\left|\left\langle C\right\rangle\right|$ obtained using the ‘Unmixed Probabilities’ method. The left half corresponds to the WMAP Q-band, and the right half to the Planck mission using the COMMANDER pipeline.
WMAP Q Band Planck COMMANDER Pipeline
133 135 137 45 2.4637 0.0009 133 135 137 45 2.2321 0.0009
133 135 139 47 2.4625 0.0009 43 45 47 135 2.2309 0.0009
133 137 139 49 2.4600 0.0009 133 135 139 47 2.2305 0.0009
Acknowledgements
This work was partially funded by the Conselleria d’Educació, Universitats i Ocupació of the Generalitat Valenciana through projects CIAICO/2022/252 and ASFAE/2022/005. We gratefully acknowledge the invaluable technical assistance of our colleague Rafael Garbayo from the Computer Administration of the Statistics Department at Universidad Miguel Hernández. We also thank Professor Arjun Berera of the University of Edinburgh for his insightful comments and bibliographic information regarding our preliminary version arXiv:2302.05125.